Find the slope of the line passing through the point (-4,-1) and the vertex of the parabola given by f(x)=x^2+2
Answer by lwsshak3(11628)

Find the slope of the line passing through the point (-4,-1) and the vertex of the parabola given by f(x)=x^2+2
vertex form of equation for a parabola:
y=(x-h)^2+k, (h,k)=(x,y) coordinates of the vertex
for given equation, y=x^2+2
vertex: (0,2)
..
equation of line:
y=mx+b, m=slope, b=y-intercept
For given line:
slope=∆y/∆x=(-1-2)/(-4-0)=-3/-4=3/4
Equation: y=3x/4+b
solve for b using coordinates of either point on the line (0,2)
2=0+b
b=2
Equation: y=3x/4+2